OpenStudy (anonymous):

When it is 2 hours after 2 o'clock, then it is 4 o'clock (2 + 2 = 4). When it is 10 hours after 10 o'clock, then it is 8 o'clock. In this kind of "clock arithmetic," 10 + 10 = 8. When a clock time gets bigger than 12, you subtract 12 and take the answer as the actual clock time. For example, if you subtract 12 from 20, the answer is 8, so 20 o'clock is really 8 o'clock.

OpenStudy (anonymous):

Brad has a certain medication that he needs to take every 5 hours without fail, starting at 1 o'clock on a certain day. The sequence of clock times that he takes his pills is 1, 6, 11, 4, 9, ... What is the clock time when Brad takes his 16th pill? Enter a number from 1 to 12.

OpenStudy (anonymous):

4

OpenStudy (anonymous):

u16=u1+(16-1)5=1+75=76 time = 76-12x6=4 so at 4

OpenStudy (anonymous):

First pill taken at 1:00 Fifteen pills left to take every 5 hours. That's 51* 5 = 75 hours after 1:00 = 76:00. 76 is congrient to 4 modulo 12; therefore, Brad takes his sixteeneh pill at 4:00.